# ModHiFi: Identifying High Fidelity predictive components for Model Modification

### v1

@@ Abstract
Open weight models, which are ubiquitous, rarely provide access to their training data or loss function. This makes modifying such models for tasks such as pruning or unlearning constrained by this unavailability an active area of research. Existing techniques typically require gradients or ground-truth labels, rendering them infeasible in settings with limited computational resources. In this work, we investigate the fundamental question of identifying components that are critical to the model’s predictive performance, without access to either gradients or the loss function, and with only distributional access such as synthetic data. We theoretically demonstrate that the global reconstruction error is linearly bounded by local reconstruction errors for Lipschitz-continuous networks such as CNNs and well-trained Transformers (which, contrary to existing literature, we find exhibit Lipschitz continuity). This motivates using the locally reconstructive behavior of component subsets to quantify their global importance, via a metric that we term Subset Fidelity. In the uncorrelated features setting, selecting individual components via their Subset Fidelity scores is optimal, which we use to propose ModHiFi, an algorithm for model modification that requires no training data or loss function access. ModHiFi-P, for structured pruning, achieves an 11% speedup over the current state of the art on ImageNet models and competitive performance on language models. ModHiFi-U, for classwise unlearning, achieves complete unlearning on CIFAR-10 without fine-tuning and demonstrates competitive performance on Swin Transformers. [^1] Our code is available at https://github.com/DhruvaKashyap/modhifi
@@ 1 Introduction
Modern deep learning has made significant strides in a wide variety of tasks, such as classification [krizhevsky2012, krizhevsky2009learning], image generation [goodfellow2020generative], and natural language processing [min2023recent]; moreover, well-trained open weight models for such tasks are easily accessible. However, significant challenges remain in their deployment, such as inference in resource-constrained settings [shuvo2022efficient, prakash2019optimizing], inference with unbalanced or biased data [hooker2019compressed, hooker2020characterising], and interpretable inference [zhang2021survey]. These challenges have increased interest in methods that modify the parameters of well-trained models to alter their behavior [santurkar2021editing, sahoo2024layerwise, shah2024decomposing]. These methods include pruning [hoefler2021sparsity], classwise unlearning [kodge2024deepunlearning, shah2024decomposing, jain2022distilling], and debiasing [shah2024decomposing, mitchell2021fast], among other model modifications. Moreover, recent work has studied model modification in the setting where the original training data and loss function are unavailable [murti2024discedit]; this is motivated by concerns related to privacy and security [yin2020dreaming], and also the use of synthetic data, which has become critical in a variety of language modeling settings [ding2024data, tan2024large]. Thus, we address the challenging problem of altering well-trained models without training data or the loss function, and only with distributional access to the original training distribution in the form of synthetic data, focusing specifically on structured pruning and classwise unlearning.
Modifying open weight models without the loss function and only synthetic data requires answering a fundamental question: which components in a model contribute significantly to its predictive performance [^2] We measure predictive performance using accuracy for classification tasks, and perplexity and other measures for language modeling tasks [shah2024decomposing, murti2024discedit].? However, most methods that identify critical components for specific modifications (e.g., pruning) cannot be applied to others (e.g., unlearning) [murti2024discedit], often require expensive fine-tuning, and are architecture-specific. Moreover, most methods utilize gradients to assess the impact of a component on the loss objective, which is not feasible in the absence of the loss function and the training data. While the LLM pruning literature uses calibration datasets to mitigate the problem of the absence of datasets [ma2023llm, ashkboos2024slicegpt], the problem of achieving sparsity in vision models without original training data is hard and unsolved [hoefler2021sparsity]. Moreover, the problem of performing classwise unlearning without original training data is unattempted [jia2023model, murti2024discedit].
@@
    Local-to-Global with Lipschitzness. An open question is the extent to which local model modifications impact the predictive performance of the model. In the absence of loss functions and training sets, estimating the impact of component modification by using gradients (as done in [liu2021group, ma2023llm, jia2023model]) is infeasible. To address this, in Section˜3.2, we show that for Lipschitz continuous networks, the reconstruction error at the final layer is at most linear in the local reconstruction errors. Moreover, contrary to the assertion that transformers are not Lipschitz continuous [qi2023lipsformer], in Corollary˜B.4, we show that this is not the case for well-trained transformers, allowing us to apply Section˜3.2 to not just CNNs, but well-trained ViTs and LLMs as well.
@@
    Identifying Subsets of Important Components. Contrary to prior work that usually infer saliencies for single components, we propose measuring the importance of sets of components for understanding cumulative effects of groups of components on a model’s predictive performance. Leveraging Section˜3.2, we propose Subset Fidelity, which quantifies the extent to which a subset of components can reconstruct the output after modifying their weights. However, computing optimal subsets is NP-complete, motivating us to compute Subset Fidelity scores for singleton sets. Section˜3.2 establishes that selecting singletons with the highest subset fidelity scores is optimal when the features are uncorrelated.
@@
    Modifying Models with ModHiFi-X. Motivated by Section˜3.2, we propose the ModHiFi algorithm, which uses the subset fidelity of singletons to modify models for pruning and classwise unlearning; the algorithm identifies important components using the singleton scores, and removes them (for classwise unlearning, ModHiFi-U) or retains them (for structured pruning, ModHiFi-P). We show that ModHiFi-P achieves state-of-the-art speedup for Imagenet models, and is consistently competitive with current baselines for language models. For classwise unlearning, ModHiFi-U achieves complete unlearning on all CIFAR-10 classes without fine-tuning, and is competitive with baselines on Swin-Transformers that finetune. When allowing for a similar fine-tuning budget as said baselines, ModHiFi-U outperforms, particularly when given access to training data. These empirical results demonstrate the practical effectiveness of Subset Fidelity.
@@ 2.1 Background and Notation
Let $p]=\{1,\ldots,p\}$ for $p\in\mathbb{N}$. We denote vectors by $\bm{v}\in\mathbb{R}^{n}$ with entries $v_{i}$, and matrices by $\bm{B}\in\mathbb{R}^{n\times m}$ with rows $\bm{b}_{i}^{\top}$ and columns $\bm{B}_{:,j}$. The vectors $\bm{1}_{d}$ and $\bm{0}_{d}$ denote the all-ones and all-zeros vectors in $\mathbb{R}^{d}$, respectively. For a vector $\bm{v}$, $\mathrm{diag}(\bm{v})$ is the corresponding diagonal matrix. We use $\|\bm{v}\|_{2}$ for the Euclidean norm and $|\bm{v}|_{0}$ for the count of nonzero entries. For matrices $\bm{C},\bm{D}$, the inner product is $\langle\bm{C},\bm{D}\rangle=\mathrm{Tr}(\bm{C}^{\top}\bm{D})$, the Frobenius norm is $\|\bm{C}\|_{F}=\sqrt{\langle\bm{C},\bm{C}\rangle}$, and the spectral norm $\|\bm{C}\|$ is the largest singular value. For index sets $A\subseteq[n]$ and $B\subseteq[m]$, $\bm{C}[A,B]$ denotes the submatrix of $\bm{C}$ defined by these indices. Expectations of a random variable $\mathrm{X}$ are written $\mathbb{E}_{\mathrm{X}}[\cdot]$, omitting the subscript when clear from context.
@@
Consider the $l$-th layer of a convolutional network. It transforms an input $\Phi^{l}(\bm{X})\in\mathbb{R}^{c_{\mathrm{in}}^{l}\times h^{l-1}\times w^{l-1}}$ (the output of preceding layers) into an output $\bm{\mathsfit{Y}}^{l}(\bm{X})\in\mathbb{R}^{c_{\mathrm{out}}^{l}\times h^{l}\times w^{l}}$. The layer is parameterized by a weight tensor $\bm{\mathsfit{W}}^{l}\in\mathbb{R}^{c_{\mathrm{out}}^{l}\times c_{\mathrm{in}}^{l}\times k^{l}\times k^{l}}$. Each output channel $c\in[c_{\mathrm{out}}^{l}]$ is computed as the sum of convolved input channels:
@@
where $\mathbin{\star}$ denotes the standard 2D convolution. We define $\bm{A}_{ci}^{l}(\bm{X})\coloneqq\Phi_{i}^{l}(\bm{X})\mathbin{\star}\bm{W}_{ci}^{l}\ \ (\in\mathbb{R}^{h^{l}\times w^{l}})$ as the input contribution from channel $i$ to output channel $c$. This decomposition is central to our component importance analysis. For notational simplicity, we omit explicit bias terms and stride/padding specifications, as our analysis generalizes to these standard configurations without loss of generality.
@@
Transformer blocks consist of Multi-Head Attention (MHA) and Feed-Forward Networks (FFN), with pre-normalization (LayerNorm or RMSNorm) [ba2016layer, ashkboos2024slicegpt]. Our analysis focuses on the FFN, treating attention as a pass-through for component importance; we leave attention-specific analysis for future work. Let the input to the $l$-th layer be $\phi^{l}(\bm{X})\in\mathbb{R}^{T\times d}$, where $T$ is sequence length and $d$ is model dimension. The FFN comprises two linear transformations, $W_{U}^{l}\in\mathbb{R}^{d\times d_{\mathrm{ff}}}$ and $W_{D}^{l}\in\mathbb{R}^{d_{\mathrm{ff}}\times d}$, and an elementwise nonlinearity $\sigma(\cdot)$ and it’s output, $\mathrm{FFN}^{l}(\phi^{l}(\bm{X}))=\sigma(\phi^{l}(\bm{X})W_{U}^{l})\,W_{D}^{l}.$ Defining the intermediate activation $\Phi^{l}(\bm{X})\coloneqq\sigma(\phi^{l}(\bm{X})W_{U}^{l})\in\mathbb{R}^{T\times d_{\mathrm{ff}}}$, the contribution from intermediate neuron $i\in[d_{\mathrm{ff}}]$ to output coordinate $c\in[d]$ is:
@@
To unify these architectures, we define a common abstraction used in our theoretical results. Let $\mathrm{N}_{\theta}=f^{L}\circ\cdots\circ f^{1}$ be a network composed of $L$ layers. Each layer $l$ maps an input $\Phi^{l}(\bm{X})$ to an output $\bm{\mathsfit{Y}}^{l}(\bm{X})$. Crucially, for both CNNs and Transformers, the output channel $c$ can be decomposed as a sum of atomic input contributions:
$\bm{Y}_{c}^{l}(\bm{X})=\sum_{i=1}^{c_{\mathrm{in}}^{l}}\bm{A}_{ci}^{l}(\bm{X}),$
where $\bm{A}_{ci}^{l}$ represents either the spatial convolution (Equation˜CONV) or the token-wise linear projection (Equation˜LIN). This additive structure allows us to analyze component fidelity in an architecture-agnostic manner.
@@ 2.2 Modifying Open Weight Models without Training Data or the Loss Function via Distributional Access
We formally define model modification as the process of selectively altering parameters of a pre-trained model, without retraining from scratch, to satisfy constraints such as efficiency, privacy, or safety [shah2024decomposing, jia2023model, hoefler2021sparsity, santurkar2021editing]. This includes tasks including structured pruning, unlearning [kurmanji2024towards], debiasing [jain2022distilling], continual or life-long learning [golkar2019continual, sahoo2024layerwise]. A major impediment in real-world modification is the unavailability of the original training data and loss function [murti2024discedit, hoefler2021sparsity]. To address this, we operate under the constraint of distributional access, specifically utilizing synthetic data [tan2024large, ding2024data], to proxy the underlying data distribution without requiring the original corpus.
@@
Problem Formulation
Let $\theta^{\star}\in\mathbb{R}^{D}$ be the parameters of a well-trained model. We seek a modification mask $\mathsf{m}^{\star}\in\mathcal{M}$ (where $\mathcal{M}$ defines permissible modifications, e.g., binary masks for pruning) to produce modified parameters $\theta^{E}=\theta^{\star}\odot\mathsf{m}^{\star}$. The optimal mask minimizes a weighted objective over distributions $\{(\mathcal{D}_{i},\alpha_{i})\}$:
@@
Structured Pruning: The goal is to maximize sparsity subject to performance retention. Here, $\mathcal{M}$ restricts masks to have at most $B$ active structured components (e.g., filters, rows). The objective is:
$\mathsf{m}^{\star}=\operatorname*{arg\,min}_{\mathsf{m}\in\mathcal{M},\|\mathsf{m}\|_{0}\leq B}\,\mathbb{E}_{\mathcal{D}}\left[\mathcal{L}(\mathrm{N}_{\theta^{\star}\odot\mathsf{m}}(\mathrm{X}))\right].$
Classwise Unlearning: The goal is to degrade performance on a forget class ($\mathcal{D}_{f}$) while maintaining performance on retain classes ($\mathcal{D}_{r}$). We formulate this as minimizing the retain loss while maximizing the forget loss:
$\mathsf{m}^{\star}=\operatorname*{arg\,min}_{\mathsf{m}\in\mathcal{M}\ (=\mathbb{R}^{D})}\left(\mathbb{E}_{\mathcal{D}_{r}}\left[\mathcal{L}(\mathrm{N}_{\theta^{\star}\odot\mathsf{m}})\right]-\mathbb{E}_{\mathcal{D}_{f}}\left[\mathcal{L}(\mathrm{N}_{\theta^{\star}\odot\mathsf{m}})\right]\right).$
Our core challenge is to solve Equation˜MODIFY using only synthetic samples, without access to the ground-truth labels or the original loss function.

### v2

@@ Abstract
Open weight models, which are ubiquitous, rarely provide access to their training data or loss function. This makes modifying such models for tasks such as pruning or unlearning, which are constrained by this unavailability, an active area of research. Existing techniques typically require gradients or ground-truth labels, rendering them infeasible in settings with limited computational resources. In this work, we investigate the fundamental question of identifying components that are critical to the model’s predictive performance, without access to either gradients or the loss function, and with only distributional access such as synthetic data. We theoretically demonstrate that the global error is linearly bounded by local reconstruction errors for Lipschitz-continuous networks such as CNNs and well-trained Transformers (which, contrary to existing literature, we find exhibit Lipschitz continuity). This motivates using the locally reconstructive behavior of component subsets to quantify their global importance, via a metric that we term Subset Fidelity. In the uncorrelated features setting, selecting individual components based on their Subset Fidelity scores is optimal, which we utilize to propose ModHiFi, an algorithm for model modification that requires neither training data nor access to a loss function. ModHiFi-P, for structured pruning, achieves an 11% speedup over the current state of the art on ImageNet models and competitive performance on language models. ModHiFi-U, for classwise unlearning, achieves complete unlearning on CIFAR-10 without fine-tuning and demonstrates competitive performance on Swin Transformers. [^1] Our code is available at https://github.com/DhruvaKashyap/modhifi
@@ 1 Introduction
Modern deep learning has made significant strides in a wide variety of tasks, such as classification [36, 37], image generation [22], and natural language processing [51]; moreover, well-trained open weight models for such tasks are easily accessible. However, significant challenges remain in their deployment, such as inference in resource-constrained settings [72, 62], inference with unbalanced or biased data [29, 30], and interpretable inference [97]. These challenges have increased interest in methods that modify the parameters of well-trained models to alter their behavior [66, 64, 68]. These methods include pruning [28], classwise unlearning [35, 68, 32], and debiasing [68, 52], among other model modifications. Moreover, recent work has studied model modification in the setting where the original training data and loss function are unavailable [55]; this is motivated by concerns related to privacy and security [92], and also the use of synthetic data, which has become critical in a variety of language modeling settings [10, 76]. Thus, we address the challenging problem of altering well-trained models without training data or the loss function, and only with distributional access to the original training distribution in the form of synthetic data, focusing specifically on structured pruning and classwise unlearning.
Modifying open weight models without the loss function and only synthetic data requires answering a fundamental question: which components in a model contribute significantly to its predictive performance [^2] We measure predictive performance using accuracy for classification tasks, and perplexity and other measures for language modeling tasks [68, 55].? However, most methods that identify critical components for specific modifications (e.g., pruning) cannot be applied to others (e.g., unlearning) [55], often require expensive fine-tuning, and are architecture-specific. Moreover, most methods utilize gradients to assess the impact of a component on the loss objective, which is not feasible in the absence of the loss function and the training data. While the LLM pruning literature uses calibration datasets to mitigate the problem of the absence of datasets [48, 2], the problem of achieving sparsity in vision models without original training data is hard and unsolved [28]. Moreover, the issue of performing classwise unlearning without access to the original training data has not been addressed [33, 55].
@@
    Local-to-Global with Lipschitzness. An open question is the extent to which local model modifications impact the predictive performance of the model. In the absence of loss functions and training sets, estimating the impact of component modification by using gradients (as done in [45, 48, 33]) is infeasible. To address this, in Section˜3.2, we show that for Lipschitz continuous networks, the reconstruction error at the final layer is at most linear in the local reconstruction errors. Moreover, contrary to the assertion that transformers are not Lipschitz continuous [63], in Corollary˜B.4, we show that this is not the case for well-trained transformers, allowing us to apply Section˜3.2 to not just CNNs, but well-trained ViTs and LLMs as well.
@@
    Identifying Subsets of Important Components. Contrary to prior work, which usually infers saliencies for single components, we propose measuring the importance of sets of components to understand the cumulative effects of groups of components on a model’s predictive performance. Leveraging Section˜3.2, we propose Subset Fidelity, which quantifies the extent to which a subset of components can reconstruct the output after modifying their weights. However, computing optimal subsets is NP-complete, motivating us to compute Subset Fidelity scores for singleton sets. Section˜3.2 establishes that selecting singletons with the highest subset fidelity scores is optimal when the features are uncorrelated.
@@
    Modifying Models with ModHiFi-X. Motivated by Section˜3.2, we propose the ModHiFi algorithm, which uses the subset fidelity of singletons to modify models for pruning and classwise unlearning; the algorithm identifies important components using the singleton scores, and removes them (for classwise unlearning, ModHiFi-U) or retains them (for structured pruning, ModHiFi-P). We demonstrate that ModHiFi-P achieves state-of-the-art speedup for ImageNet models and consistently competes with current baselines for language models. For classwise unlearning, ModHiFi-U achieves complete unlearning on all CIFAR-10 classes without fine-tuning and is competitive with baselines on Swin-Transformers that require fine-tuning. When allowing for a similar fine-tuning budget as said baselines, ModHiFi-U outperforms, particularly when given access to training data. These empirical results demonstrate the practical effectiveness of Subset Fidelity.
@@ 2.1 Background and Notation
Let $p]=\{1,\ldots,p\}$ for $p\in\mathbb{N}$. We denote vectors by $\bm{v}\in\mathbb{R}^{n}$ with entries $v_{i}$, and matrices by $\bm{B}\in\mathbb{R}^{n\times m}$ with rows $\bm{b}_{i}^{\top}$ and columns $\bm{B}_{:,j}$. The vectors $\bm{1}_{d}$ and $\bm{0}_{d}$ denote the all-ones and all-zeros vectors in $\mathbb{R}^{d}$, respectively. We use $\|\bm{v}\|_{2}$ for the Euclidean norm. For matrices $\bm{C},\bm{D}$, the inner product is $\langle\bm{C},\bm{D}\rangle=\mathrm{Tr}(\bm{C}^{\top}\bm{D})$, the Frobenius norm is $\|\bm{C}\|=\sqrt{\langle\bm{C},\bm{C}\rangle}$, and the spectral norm $\|\bm{C}\|_{2}$ is the largest singular value. For index sets $A\subseteq[n]$ and $B\subseteq[m]$, $\bm{C}[A,B]$ denotes the submatrix of $\bm{C}$ defined by these indices. Expectations of a random variable $\mathrm{X}$ are written $\mathbb{E}_{\mathrm{X}}[\cdot]$, omitting the subscript when clear from context.
@@
Consider the $l$-th layer of a convolutional network. It transforms an input (the output of preceding layers) $\Phi^{l}(\bm{X})\in\mathbb{R}^{c_{\mathrm{in}}^{l}\times h^{l-1}\times w^{l-1}}$ into output $\bm{\mathsfit{Y}}^{l}(\bm{X})\in\mathbb{R}^{c_{\mathrm{out}}^{l}\times h^{l}\times w^{l}}$. The layer is parameterized by a weight tensor $\bm{\mathsfit{W}}^{l}\in\mathbb{R}^{c_{\mathrm{out}}^{l}\times c_{\mathrm{in}}^{l}\times k^{l}\times k^{l}}$. Each output channel $c\in[c_{\mathrm{out}}^{l}]$ is computed as the sum of convolved input channels:
@@
where $\mathbin{\star}$ denotes the standard 2D convolution. We define $\bm{A}_{ci}^{l}(\bm{X})\coloneqq\Phi_{i}^{l}(\bm{X})\mathbin{\star}\bm{W}_{ci}^{l}\ \ (\in\mathbb{R}^{h^{l}\times w^{l}})$ as the input contribution from channel $i$ to output channel $c$. For notational simplicity, we omit explicit bias terms and stride/padding specifications, as our analysis generalizes to these standard configurations without loss of generality.
@@
Transformer blocks consist of Multi-Head Attention (MHA) and Feed-Forward Networks (FFN), with pre-normalization (LayerNorm or RMSNorm) [3, 2]. Our analysis focuses on the FFN; we leave attention-specific analysis for future work. Let the input to the $l$-th layer be $\phi^{l}(\bm{X})\in\mathbb{R}^{T\times d}$, where $T$ is sequence length and $d$ is model dimension. The FFN comprises two linear transformations, $W_{U}^{l}\in\mathbb{R}^{d\times d_{\mathrm{ff}}}$ and $W_{D}^{l}\in\mathbb{R}^{d_{\mathrm{ff}}\times d}$, and an elementwise nonlinearity $\sigma(\cdot)$ and it’s output, $\mathrm{FFN}^{l}(\phi^{l}(\bm{X}))=\sigma(\phi^{l}(\bm{X})W_{U}^{l})\,W_{D}^{l}.$ Defining the intermediate activation $\Phi^{l}(\bm{X})\coloneqq\sigma(\phi^{l}(\bm{X})W_{U}^{l})\in\mathbb{R}^{T\times d_{\mathrm{ff}}}$, the contribution from intermediate neuron $i\in[d_{\mathrm{ff}}]$ to output coordinate $c\in[d]$ is:
@@
To unify these architectures, we define a common abstraction used in our theoretical results. Let $\mathrm{N}_{\theta}=f^{L}\circ\cdots\circ f^{1}$ be a network composed of $L$ layers. Each layer $l$ maps an input $\Phi^{l}(\bm{X})$ to an output $\bm{\mathsfit{Y}}^{l}(\bm{X})$. Crucially, for both CNNs and Transformers, the output channel $c$ can be decomposed as a sum of atomic input contributions: $\bm{Y}_{c}^{l}(\bm{X})=\sum_{i=1}^{c_{\mathrm{in}}^{l}}\bm{A}_{ci}^{l}(\bm{X}),$ where $\bm{A}_{ci}^{l}$ represents either the spatial convolution (Eq.˜CONV) or the token-wise linear projection (Eq.˜LIN). This decomposition is central to our analysis of component importance. This additive structure allows us to analyze component fidelity in an architecture-agnostic manner.
@@ 2.2 Modifying Open Weight Models without Training Data or the Loss Function via Distributional Access
We formally define model modification as the process of selectively altering parameters of a pre-trained model, without retraining from scratch, to satisfy constraints such as efficiency, privacy, or safety [68, 33, 28, 66]. This includes tasks including structured pruning, unlearning [38], debiasing [32], continual or life-long learning [21, 64]. A major impediment in real-world modification is the unavailability of the original training data and loss function [55, 28]. To address this, we operate under the constraint of distributional access, specifically utilizing synthetic data [76, 10], to proxy the underlying data distribution without requiring the original corpus.
@@
Formulating Model Modification
Let $\theta^{\star}\in\mathbb{R}^{D}$ be the parameters of a well-trained model. We seek a modification mask $\mathsf{m}^{\star}\in\mathcal{M}$ (where $\mathcal{M}$ defines permissible modifications, e.g., binary masks for pruning) to produce modified parameters $\theta^{E}=\theta^{\star}\odot\mathsf{m}^{\star}$. Given data distributions $\{\mathcal{D}_{i}\}_{i=1}^{K}$ and weights $\bm{\alpha}\in\mathbb{R}^{K}$, the optimal modification mask is defined as:
@@
Structured Pruning The goal is to maximize performance subject to sparsity. Let the parameters be partitioned into $G$ disjoint structured groups $\{\mathcal{G}_{g}\}_{g=1}^{G}$ (e.g., filters, channels, rows, or columns), with $\theta^{\star}=(\theta^{\star}_{\mathcal{G}_{1}},\dots,\theta^{\star}_{\mathcal{G}_{G}})$. The admissible set enforces a sparsity budget $B$, $\mathcal{M}_{\mathrm{SP}}:=\{\mathsf{m}\in\mathbb{R}^{D}~|~\exists\,\mathsf{z}\in\{0,1\}^{G},\;\bm{\delta}\in\mathbb{R}^{D}\text{ s.t. }\mathsf{m}_{j}=\mathsf{z}_{g}\delta_{j}\;\forall j\in\mathcal{G}_{g},\;\sum_{g=1}^{G}z_{g}\leq B\}$. Structured pruning is recovered from (MODIFY) by setting $K=1$, $\mathcal{D}_{1}=\mathcal{D}$ (original task distribution), $\alpha_{1}=1$, and $\mathcal{M}=\mathcal{M}_{\mathrm{SP}}$, yielding
$\mathsf{m}^{\star}=\operatorname*{arg\,min}_{\mathsf{m}\in\mathcal{M}_{SP}}\,\mathbb{E}_{\mathcal{D}}\left[\mathcal{L}(\mathrm{N}_{\theta^{\star}\odot\mathsf{m}}(\mathrm{X}))\right].$ (STRUCT-PRUNE)
Classwise Unlearning The goal is to degrade performance on a forget distribution $\mathcal{D}_{f}$ while preserving performance on a retain distribution $\mathcal{D}_{r}$. We impose no additional structural constraints on the modification and set $\mathcal{M}_{U}=\mathbb{R}^{D}$. Classwise unlearning is obtained from (MODIFY) by setting $K=2$, $(\mathcal{D}_{1},\mathcal{D}_{2})=(\mathcal{D}_{r},\mathcal{D}_{f})$, $(\alpha_{1},\alpha_{2})=(1,-1)$, and $\mathcal{M}=\mathcal{M}_{U}$, yielding
$\mathsf{m}^{\star}~=~\operatorname*{arg\,min}_{\mathsf{m}\in\mathcal{M}_{U}}\mathbb{E}_{\mathrm{X}\sim\mathcal{D}_{r}}[\mathcal{L}(\mathrm{N}_{\theta^{\star}\odot\mathsf{m}}(\mathrm{X}))]-\mathbb{E}_{\mathrm{X}\sim\mathcal{D}_{f}}[\mathcal{L}(\mathrm{N}_{\theta^{\star}\odot\mathsf{m}}(\mathrm{X}))].$ (UNLEARN)
Our core challenge is to solve (MODIFY) using only synthetic samples, without access to ground-truth labels or the original loss.